If a Home Depot worker can paint a house in 4 hours, and a Lowes worker can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
Answer by stanbon(75887)

If a Home Depot worker can paint a house in 4 hours, and a Lowes worker can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
-------------------
1st worker DATA;
Time = 4 hrs/job ; rate = 1/4 job/hr
--------------
2nd workier DATA:
Time = 6 hr/job ; rate = 1/6 job/hr
----------------
Together DATA:
time = x hr/job ; rate = 1/x job/hr
====================
EQUATION:
rate + rate = together rate
1/4 + 1/6 = 1/x
3x+ 2x = 12
5x = 12
x = 2.4 hrs (time for them to do the job together
======================
Cheers,
Stan H.